The arrival rate of customers arriving at a bank counter follows a Poisson distribution with a mean rate of 4 per 10 minutes interval. Find the probability that i. Exactly 0 customer will arrive in 10 minutes interval ii. Exactly 2 customers will arrive in 10 minutes interval iii. At most 2 customers will arrive in 10 minutes interval iv. At least 3 customers will arrive in 10 minutes interval 


1
Expert's answer
2021-04-14T11:50:32-0400

Let "N" be the number of customers arriving in 10 minutes interval.

"N" has "Poisson(4)" distribution, hence "P(N=n)=\frac{4^ne^{-4}}{n!}".

i. "P(N=0)=\frac{4^0e^{-4}}{0!}=e^{-4}".

ii. "P(N=2)=\frac{4^2e^{-4}}{2!}=8e^{-4}".

iii.

"P(N\leq2)=\\
P(N=0)+P(N=1)+P(N=2)=\\
e^{-4}+\frac{4^1e^{-4}}{1!}+8e^{-4}=13e^{-4}."

iv.

"P(N\geq3)=P(N>2)=1-P(N\leq2)=1-13e^{-4}."

Answer: i. "e^{-4}\approx0.0183", ii. "8e^{-4}\approx 0.1465", iii. "10e^{-4}\approx 0.1831", iv. "1-13e^{-4}\approx 0.7619".